OpenStudy (anonymous):

write the quadratic function in standard form that has its vertex at 3,-2 and passes through 1,6

OpenStudy (anonymous):

what's the vertex?

OpenStudy (anonymous):

3,-2*

OpenStudy (campbell_st):

ok... so the vertex form of a quadratic is \[y = a(x -h)^2 + k\] you know h = 3 and k = -2 substitute then to get the equation next you need to find a substitute x = 1 and y = 6 and solve for a

jimthompson5910 (jim_thompson5910):

Vertex form: y = a(x-h)^2 + k we know the vertex is (3, -2), so (h,k) = (3, -2) ---> h = 3 and k = -2 y = a(x-h)^2 + k y = a(x-3)^2 + (-2) y = a(x-3)^2 - 2 Now plug in (x,y) = (1,6) or x = 1 and y = 6 to get y = a(x-3)^2 - 2 6 = a(1-3)^2 - 2 6 = a(-2)^2 - 2 6 = 4a - 2 keep going to solve for 'a'

OpenStudy (mertsj):